OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle whose legs are 3 and √2

OpenStudy (anonymous):

Take the square of both legs and add them together. (3^2 + (sqrt(2))^2 = 9+2 = 11) THis is the square of the length of the hypotenuese.

OpenStudy (anonymous):

So the length of the hypotenuse is \[\sqrt{11}\]